OpenStudy (anonymous):

Solving quadratic Equations by factoring: 8x^2+2x=3

OpenStudy (141823):

do you have to use quadratic formula or any method?

OpenStudy (anonymous):

Quadratic Formula

OpenStudy (anonymous):

What??

OpenStudy (kidrah69):

\[\frac{ x = -(b) ± \sqrt b^2 - 4(a)(c)}{ 2(a)}\]

OpenStudy (kidrah69):

We want to get a C value too so subtract 3 from both sides and you get \[8x^2+2x-3=0\]

OpenStudy (kidrah69):

So: A=8 B=2 C=-3

OpenStudy (anonymous):

Then you factor it out right?

OpenStudy (kidrah69):

yep

OpenStudy (anonymous):

and you get (4x-1)(2x-3)

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

And for your zeros you get 1/4 and -6/2 I think?

OpenStudy (kidrah69):

lets do it step by step \[\frac{ x=−(2)±√2^2−4(8)(-3) }{ 2(8) }\] now when you do this you get... \[\frac{ -2 ±\sqrt{100}}{ 16 }\]

OpenStudy (anonymous):

Whoa.. I have never done it like that. Im only in algebra one?

OpenStudy (kidrah69):

so am i :) this is an easier way

OpenStudy (anonymous):

Oh okay

OpenStudy (kidrah69):

Do you get it uptil this part?

OpenStudy (anonymous):

Yes your just simplifying, right?

OpenStudy (kidrah69):

Now all you do is split the equation :) \[\frac{ -2+\sqrt{100} }{ 16 }\] and \[\frac{ -2-\sqrt{100} }{ 16 }\]

OpenStudy (anonymous):

And you get your zeros, So you get 1/2 and -3/4????

OpenStudy (kidrah69):

1)-3/4 2)1/2 yep :)

OpenStudy (anonymous):

Wow thanks! Can I do a few more with you?

OpenStudy (mathmath333):

is factoring and using quadratic formula same

OpenStudy (anonymous):

Just to make sure I get it?

OpenStudy (kidrah69):

Yeah but i need to make sure i've done this right :) @mathmath333 @perl is it correct?

OpenStudy (anonymous):

Okay well I'll put the equation in: So the equation is 3x^2-x=4

OpenStudy (kidrah69):

Can you make another post for it please

OpenStudy (mathmath333):

looks like answer is correct

OpenStudy (kidrah69):

Yay!!! :D

OpenStudy (kidrah69):

Oh you wanted it by factoring :/ i dont know that way i only know this way

OpenStudy (anonymous):

Whats this way?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} &8x^2+2x=3\hspace{.33em}\\~\\ &\implies 8x^2+2x-3=0\hspace{.33em}\\~\\ &\implies \color{red}{8}x^2+\color{blue}{6}x-\color{blue}{4}x-\color{red}{3}=0\hspace{.33em}\\~\\ &\implies 2x(4x+3)-(4x+3)=0\hspace{.33em}\\~\\ &\implies (2x-1)(4x+3)=0\hspace{.33em}\\~\\ \end{align}}\) note that \(\large \color{black}{\begin{align} \color{red}{8\times 3}=\color{blue}{6\times 4} \end{align}}\)

OpenStudy (kidrah69):

this way was simplifying (the one i did)

OpenStudy (mathmath333):

u have to write \(2x\) (here \(6x-4x\)) such that it equals \(=8\times 3=24\) and \(6\times 4=24\)

OpenStudy (anonymous):

Oh does it still give the same answer I get when factoring and thank you @mathmath333

OpenStudy (mathmath333):

yes it gives same answer solve further and check \(\large \color{black}{\begin{align} \ &\implies (2x-1)(4x+3)=0\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

1/2 and -3/4??

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

Awesome! Thanks!